# Second-Order Constrained Dynamic Optimization

### v2

@@ 3 Unconstrained Differential Dynamic Programming
where $\delta x_{k}:=x_{k}-\bar{x}_{k}$, $\delta u_{k}:=u_{k}-\bar{u}_{k}$ are deviations from the nominal sequences. The RHS is expanded in the same way about $\bar{x}_{k}$, $\bar{u}_{k}$, and $\bar{x}_{k+1}$, where $\delta\bar{x}_{k+1}$ is eliminated by plugging in linearized dynamics $\delta x_{k+1}=f_{x}\delta x_{k}+f_{u}\delta u_{k}$. Mapping the terms of both sides of the expanded (32) gives
${Q}_{{xx},k}={l}_{{xx}}+{f_{x}}^{\mathsf{T}}{V}_{{xx},k+1}{f_{x}},\hskip 2.27621pt{Q}_{{x},k}={l}_{{x}}+{f_{x}}^{\mathsf{T}}{V}_{{x},k+1},$
${Q}_{{uu},k}={l_{{uu}}}+{f_{u}}^{\mathsf{T}}{{V}_{{xx},k+1}}{f_{u}},\hskip 2.27621pt{Q}_{{u},k}={l_{u}}+{f_{u}}^{\mathsf{T}}{{V}_{{x},k+1}},$
${Q}_{{xu},k}={l_{{xu}}}+{f_{x}}^{\mathsf{T}}{{V}_{{xx},k+1}}{f_{u}}.$
After plugging (32) and (33) into (31), we can explicitly optimize the value function with respect to $\delta{u}_{k}$ by taking a partial derivative and setting it to zero, obtaining the locally optimal control update
@@
where $\kappa_{k}$ and $K_{k}$ are known as feedforward and feedback gains, respectively. Note that we have dropped the time indices for $Q$ to lighten the notation. To ensure convergence, $Q_{{uu}}$ must be regularized when it is not PD (Yakowitz and Shoemaker, 1991), which is achieved with
@@ 4.1 Log Barrier DDP
where the second term of $\hat{Q}_{uu}$ can be omitted for the GN approximation as in section 2.2. This approximation was rederived by augmenting the barrier term as an additional element of the state in dynamics in Almubarak et al. (2022). In this work, only the first-order derivatives of constraints contribute to the second-order derivatives of $Q$. The approximation is also indirectly regularizing the Hessian, making the optimization problem more well-conditioned for DDP. There exist techniques that relax the $\log$ barrier to facilitate optimization, accepting constraint violation Grandia et al. (2019). In this paper, however, we use the exact (not relaxed) barrier function because it can keep trajectories always feasible even before convergence. This is a unique algorithm property that distinguishes the method from others. As mentioned in section 2.2, $\mu$ needs to be reduced as optimization progresses. However, as reported in previous work, a single small value of $\mu$ is sufficient in most cases.
@@ 4.3 Interior Point DDP
for $k=1\cdots N-1$. The step size for $\lambda$ is obtained without the line search because it does not affect the filter. Unfortunately, applying this rule to DDP is not as straightforward as applying a similar rule to IP for static problems due to the existence of feedback terms in DDP. Because search directions are defined via the linear feedback equations in DDP as (4.3), $\alpha_{\lambda}$ is affected by $\alpha_{s}$ through $\delta x_{k}$. In the original work, without setting the max. values of each $\alpha$ s, the line search is performed with a common parameter for $\alpha_{s}$ and $\alpha_{\lambda}$, making the line search unnecessarily conservative. This is because each step size is affected by that of the paired variable. To alleviate this problem, we propose applying the rule with different $\alpha$ for $s$ and $\lambda$ as in the original formulation of the IP method. We first find $\alpha_{s}^{\rm{max}}$ using dynamics. Since there is no closed-form solution for $\alpha_{s}^{\rm{max}}$, we rely on a line search to find $\alpha_{s}^{\rm{max}}$. Next, the line search filter is performed, setting $\alpha_{s}^{\rm{max}}$ as the upper bound. Inside of this line search, the inner line search for $\alpha_{\lambda}$ is also performed, using $\delta x_{k}$ generated with $\alpha_{s}$ from the outer line search. The IP DDP algorithm used in section 7 is implemented with this modified line search filter.
4.4 Primal Dual Augmented Lagrangian DDP
@@ 4.4.1 Single-shooting
$=$ $\min_{U,\Lambda}\sum_{k=1}^{N-1}\left[l(x_{k},u_{k})+\lambda_{\rm{e},k}^{\mathsf{T}}\Big{[}[G^{\dagger}_{k}]_{+}-P_{k}^{-1}\Big{(}\lambda_{\rm{e},k}-\frac{\lambda_{k}}{2}\Big{)}\right]$
$\ \left.+\sum_{i=1}^{w}\frac{\rho_{k,i}}{2}\left\lVert\big{[}G^{\dagger}_{k,i}\big{]}_{+}-\frac{1}{\rho_{k,i}}\Big{(}[\lambda_{\rm{e}}]_{k,i}-\frac{\lambda_{k,i}}{2}\Big{)}\right\rVert^{2}\right.$
@@
$\text{with}\ G^{\dagger}_{k}=g(x_{k},u_{k})+P_{k}^{-1}(\lambda_{\rm{e},k}-{\lambda_{k}}/{2}),$
@@
$\hat{Q}_{y}=Q_{y}+[g_{y}]_{\mathcal{A}}^{\mathsf{T}}[2\pi_{\rm{I}}-\lambda_{\rm{e}}]_{+},$
$\hat{Q}_{\lambda}=-\frac{1}{2}P^{-1}\big{[}[2\pi_{\rm{I}}-\lambda_{\rm{e}}]_{+}-\lambda_{k}\big{]},\ \hat{Q}_{\lambda y}=-[g_{y}]_{\mathcal{A}},$
$\hat{Q}_{yy}=Q_{yy}+[2\pi_{\rm{I}}-\lambda_{\rm{e}}]_{+}[g_{yy}]_{\mathcal{A}}+2[g_{y}]^{\mathsf{T}}_{\mathcal{A}}P[g_{y}]_{\mathcal{A}},$ (52)
@@
and projection in (20). $[I_{w}]_{\mathcal{A}}$ is a modified identity matrix whose $i$ th diagonal element is 1 if $g_{i}$ is active (positive after projection in (20)) and zero otherwise. First order optimality condition for quadratic approximation of $\hat{Q}$ gives
@@ 4.4.2 Multiple-shooting
$\hat{Q}_{\tilde{y}}=l_{\tilde{y}}+h_{\tilde{y}}^{\mathsf{T}}[2\pi_{\rm{E}}-\nu_{\rm{e}}]+[g_{\tilde{y}}]_{\mathcal{A}}^{\mathsf{T}}[2\pi_{\rm{I}}-\lambda_{\rm{e}}]_{+},$
$\hat{Q}_{\lambda}=-\frac{1}{2}P_{\rm{I}}^{-1}\big{[}[2\pi_{\rm{I}}-\lambda_{\rm{e}}]_{+}-\lambda_{k}\big{]},\ \hat{Q}_{\nu}=-P_{\rm{E}}^{-1}[\pi_{\rm{E}}-\nu_{\rm{e}}],$
@@ 4.5.1 Introducing Copy Variables
Here, we use $Y=[X^{\mathsf{T}},U^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{n+m}$, which is similar to $y_{k}$. This formulation is equivalent to setting $x=[X,U]$, $z=[X^{\mathrm{c}},U^{\mathrm{c}}]$, $A=I,B=-I$, and $c=0$ in (22). The AL of the problem is given by
@@ 4.5.2 ADMM DDP
ADMM DDP repeats the three update processes in (54), (57), and (58) until the residuals mentioned in section 2.6 become small enough. Upon convergence, the original and copy variables will reach to consensus, and as a result, the final solution will be optimal while satisfying all constraints. It is well known however, that ADMM might require many iterations until reaching to high accuracy Boyd et al. (2011).
@@ 5.2 Multiple shooting-SQP
where Lagrangian is given by $\mathcal{L}=J+\lambda_{\rm{v}}^{\mathsf{T}}G(X,U)+\nu_{\rm{v}}^{\mathsf{T}}\hat{F}(X,U),$ with Lagrangian multipliers $\lambda_{\rm{v}}\in\mathbb{R}^{wN}$ and $\nu_{\rm{v}}\in\mathbb{R}^{nN}$. After solving the problem above, $Y$ (both $X$ and $U$) is updated by the following linear update law:
@@ 6 Theoretical Time Complexity
In SQPs, the bottleneck is solving the QP subproblems, whose Hessian has the size of $(N-1)m\times(N-1)m$ in single- and $\{(N-1)m+Nn\}\times\{(N-1)m+Nn\}$ in multiple-shooting. These large Hessians lead to $\mathcal{O}(N^{3}m^{3})$ and $\mathcal{O}(N^{3}(m+n)^{3})$ for single- and multiple-shooting, respectively. However, as mentioned in section 5.2, several methods in multiple-shooting SQP that can reduce the complexity to $\mathcal{O}(N(m+n)^{3})$ are available. From a computational point of view, the single-shooting variant is useful for a problem with a short time horizon $N$ and a large state dimension $n$.
@@ 7 Numerical experiments
    How robust is each algorithm to varying initial conditions?
@@
To answer these three questions, we compare the constrained DDPs and SQPs on three different dynamical systems, that is, an inverted pendulum, a 2D quadrotor with a pendulum (quadpend) based on Singh et al. (2022), and a tadpole-like swimmer as in Tassa et al. (2007). For investigating (R1), we use all systems, while for (R2) and (R3), we focus on the quadpend. The state $x$ of the inverted pendulum consists of the angle and angular velocity of the pendulum as $x=[\theta,\dot{\theta}]^{\mathsf{T}}\in\mathbb{R}^{2}$ and the control $u\in\mathbb{R}$ is the torque applied to the pendulum. The state of the quadpend consists of the 2D position and orientation of the quadrotor, the angle of the pendulum, and the time derivative of them, which leads to $x\in\mathbb{R}^{8}$. We control the force that the two rotors generate, and therefore $u\in\mathbb{R}^{2}$. We choose this example because of the nonlinearity of the constraints by the pendulum part. The swimmer is a more complex system consisting of five links, whose control is torque is applied in the four joints. The dynamics of the swimmer are found in Tassa et al. (2007). The state consists of the 2D position of the nose, four joint angles and their time derivatives. Hence, the state is $x\in\mathbb{R}^{12}$. The control $u\in\mathbb{R}^{4}$ is the torques generated in the joints. Details of the dynamics of the systems and parameters are provided in Appendix D. All experiments were performed with MATLAB (MATLAB, 2022).
@@ 7.1.1 Progress of Optimization
as a representative of the gradient of cost of a trajectory. In a constrained setting, the gradient of $J$ changes to a gradient of the modified objective depending on the algorithm, as shown in the left column of Table 2. For SQP, we use the gradient of the Lagrangian in (26). We also monitor the values presented in the right column of the table, all of which are related to constraint violation, the residual of optimality condition for dual variables (if the method takes them into account), and penalty parameters (if the method includes any and changes them over iterations). The norm here is the infinity norm taken over all the time steps, e.g.,
@@ 7.1.2 Results
Inverted pendulum: In this task, the goal is to swing up the pendulum while satisfying the constraints $-0.8\leq u_{k}\leq 0.8,-1.5\leq\dot{\theta}\leq 1.5$. The results are provided in Fig. 1a for single-shooting and Fig. 1b for multiple-shooting methods. In the right column of each figure, we present the progress of optimization with the gradient of the cost and the other metrics given in Table 2. The vertical dotted lines in AL, PDAL, and ADMM DDP represent different outer loops. In the left column, graphs showing the evolution of $\theta$, $\dot{\theta}$ and control $u$ are illustrated. The constraints and targets are given as dotted lines and circles, respectively. Since constraints are linear in this problem, the Hessian of the constraints is zero. Thus, the exact Hessian and that with the GN approximation are identical. See AL multi exact and AL multi approx. in Fig. (1b). $\log$ barrier and IP DDP terminate earlier than other methods because the descent direction cannot be found with a large regularizer in (35). Fig. 2a shows the evolution of cost (original cost $J$ in (38)). The largest constraint violation (if any) over iterations is shown in Fig. 2b. We observe that most methods are able to complete the task with low $(<10^{-5})$ constraint violation. In multiple-shooting DDPs, inequality and equality constraints exhibit comparable magnitudes, contrasting with multiple-shooting SQP, where inequality constraints consistently maintain feasibility. We postulate that this phenomenon arises from the property of SQP, that is, solving QP under constraints, rather than incorporating all elements into the $Q$ function as in DDP. Although both SQP and DDP rely on line search to find a proper step size, SQP seems to better capture the information of constraints. The $\log$ barrier method can keep the trajectory always feasible and therefore does not appear in Fig. 2b. However, its cost is higher than those of the other DDP methods. This is due to the fixed penalty parameter $\mu$ in (4.1). In exchange for a simple implementation, this fixed $\mu$ makes the algorithm only approximately solve the problem, resulting in a higher cost in this experiment. In this problem, the controller needs to hit its limit over many time steps to complete the task. This makes the problem difficult for ADMM DDP, as we analyzed in section 4.6. In Appendix D.5, we further relax the control limit to $u\in[-0.9,0.9]$ and observe that ADMM DDP is able to handle problems with less tight constraints. For SQP, the multiple shooting version performs similarly to other successful DDP methods, whereas the single shooting variant performs poorly. This difference comes from the fact that the cost function in the multiple-shooting SQP is strictly quadratic in the decision variables (state and control); on the other hand, in single-shooting, the cost is not exactly quadratic in control due to the elimination of state via linearized dynamics.
@@
Quadpend: The quadpend is navigated to reach the target $x_{\rm{g}}=[2.5,-1,0,{\pi}/{2},0,0,0,0]^{\mathsf{T}}$ (pendulum upright) from the initial state $x_{0}=[-2,1,0,0,0,0,0,0]^{\mathsf{T}}$ (pendulum down) while avoiding four obstacles under box control constraints. The initial trajectory is shown in Fig. 4a, which is obtained with single-shooting PDAL DDP. We show the results of single and multiple shooting methods in Fig. 3a and Fig. 3b, respectively. In this experiment, multiple-shooting AL DDPs reached maximum iteration by reaching a plateau where the inner loops could not find a descent direction. The multiple-shooting SQP algorithm terminated due to failure in the line search to find a descent direction even after regularizing the Hessian as explained in Section 5.1. Here, the GN approximation makes a difference because of nonlinear constraints. Indeed, we observe the effectiveness of the approximation as multiple-shooting AL DDP fails without the approximation, getting stuck at a poor local minimum with a large constraint violation. Other than the multiple-shooting AL DDP with exact Hessian, all algorithms can let the quadpend hit the target with the pendulum up, although the solutions vary. This is because the algorithms are local methods that use the local approximation of the cost and dynamics.

### v3

@@ 3 Unconstrained Differential Dynamic Programming
where $\delta x_{k}:=x_{k}-\bar{x}_{k}$, $\delta u_{k}:=u_{k}-\bar{u}_{k}$ are deviations from the nominal sequences. The RHS is expanded in the same way about $\bar{x}_{k}$, $\bar{u}_{k}$, and $\bar{x}_{k+1}$, where $\delta\bar{x}_{k+1}$ is eliminated by plugging in quadratic approximation of dynamics:
$\delta x_{k+1}\approx f(\bar{x}_{k},\bar{u}_{k})+{f}_{{x}}\delta{x}_{k}+{f_{u}}\delta{u}_{k}$
$\ +\textstyle{\frac{1}{2}}({\delta{x}}_{k}^{\mathsf{T}}{f}_{{xx}}\delta{x}_{k}+2{\delta{x}}_{k}^{\mathsf{T}}{f}_{{xu}}\delta{u}_{k}+{\delta{u}}_{k}^{\mathsf{T}}{f}_{uu}\delta{u}_{k}).$
Mapping the terms of both sides of the expanded (32) gives
${Q}_{{x},k}={l}_{{x}}+{f_{x}}^{\mathsf{T}}{V}_{{x},k+1},\ \ {Q}_{{u},k}={l_{u}}+{f_{u}}^{\mathsf{T}}{{V}_{{x},k+1}},$
${Q}_{{xx},k}={l}_{{xx}}+{f_{x}}^{\mathsf{T}}{V}_{{xx},k+1}{f_{x}}+V_{x,k+1}\cdot f_{xx},$
${Q}_{{uu},k}={l_{{uu}}}+{f_{u}}^{\mathsf{T}}{{V}_{{xx},k+1}}{f_{u}}+V_{x,k+1}\cdot f_{uu},$
${Q}_{{xu},k}={l_{{xu}}}+{f_{x}}^{\mathsf{T}}{{V}_{{xx},k+1}}{f_{u}}+V_{x,k+1}\cdot f_{xu},$
where $\cdot$ in the second-order terms stands for tensor contraction along the first dimension. After plugging (32) and (33) into (31), we can explicitly optimize the value function with respect to $\delta{u}_{k}$ by taking a partial derivative and setting it to zero, obtaining the locally optimal control update
@@
where $\kappa_{k}$ and $K_{k}$ are known as feedforward and feedback gains, respectively. Note that we have dropped the time indices for $Q$ to lighten the notation. To ensure convergence, $Q_{{uu}}$ must be regularized when it is not PD (Liao and Shoemaker, 1991), which is achieved with
@@ 4.1 Log Barrier DDP
where the second term of $\hat{Q}_{uu}$ can be omitted for the GN approximation as in section 2.2. This approximation was rederived by augmenting the barrier term as an additional element of the state in dynamics in Almubarak et al. (2022). In this work, only the first-order derivatives of constraints contribute to the second-order derivatives of $Q$. The approximation is also indirectly regularizing the Hessian, making the optimization problem more well-conditioned for DDP. There exist techniques that relax the $\log$ barrier to facilitate optimization, accepting constraint violation Grandia et al. (2019). In this paper, however, we use the exact (not relaxed) barrier function because it can keep trajectories always feasible even before convergence. This is a unique algorithm property that distinguishes the method from others. As mentioned in section 2.2, $\mu$ needs to be reduced as optimization progresses. However, as reported in previous work, a single small value of $\mu$ is sufficient in most cases Almubarak et al. (2022).
@@ 4.3 Interior Point DDP
for $k=1\cdots N-1$. The step size for $\lambda$ is obtained without the line search because it does not affect the filter. Unfortunately, applying this rule to DDP is not as straightforward as applying a similar rule to IP for static problems due to the existence of feedback terms in DDP. Because search directions are defined via the linear feedback equations in DDP as (4.3), $\alpha_{\lambda}$ is affected by $\alpha_{s}$ through $\delta x_{k}$. In the original work, without setting the max. values of each $\alpha$ s, the line search is performed with a common parameter for $\alpha_{s}$ and $\alpha_{\lambda}$, making the line search unnecessarily conservative. This is because each step size is affected by that of the paired variable. To alleviate this problem, we propose applying the rule with different $\alpha$ for $s$ and $\lambda$ as in the original formulation of the IP method. We first find $\alpha_{s}^{\rm{max}}$ using dynamics. Since there is no closed-form solution for $\alpha_{s}^{\rm{max}}$, we rely on a line search to find $\alpha_{s}^{\rm{max}}$. Next, the line search filter is performed, setting $\alpha_{s}^{\rm{max}}$ as the upper bound. Inside of this line search, the inner line search for $\alpha_{\lambda}$ is also performed, using $\delta x_{k}$ generated with $\alpha_{s}$ from the outer line search. The IP DDP algorithm used in section 7 is implemented with this modified line search filter. Finally, we note that the feasibility restoration mechanism explained in 2.4 is not implemented with DDP.
4.4 Primal-Dual Augmented Lagrangian DDP
@@ 4.4.1 Single-shooting
$=$ $\min_{U,\Lambda}\sum_{k=1}^{N-1}\left[l(x_{k},u_{k})+\lambda_{\mathrm{e},k}^{\mathsf{T}}\Big{[}[G^{\dagger}_{k}]_{+}-P_{k}^{-1}\Big{(}\lambda_{\mathrm{e},k}-\frac{\lambda_{k}}{2}\Big{)}\right]$
$\ \left.+\sum_{i=1}^{w}\frac{\rho_{k,i}}{2}\left\lVert\big{[}G^{\dagger}_{k,i}\big{]}_{+}-\frac{1}{\rho_{k,i}}\Big{(}[\lambda_{\mathrm{e},k}]_{i}-\frac{\lambda_{k,i}}{2}\Big{)}\right\rVert^{2}\right.$
@@
$\text{with}\ G^{\dagger}_{k}=g(x_{k},u_{k})+P_{k}^{-1}(\lambda_{\mathrm{e},k}-{\lambda_{k}}/{2}),$
@@
$\hat{Q}_{y}=Q_{y}+[g_{y}]_{\mathcal{A}}^{\mathsf{T}}[2\pi_{\rm{I}}-\lambda_{k}]_{+},$
$\hat{Q}_{\lambda}=-\frac{1}{2}P^{-1}\big{[}[2\pi_{\rm{I}}-\lambda_{k}]_{+}-\lambda_{k}\big{]},\ \hat{Q}_{\lambda y}=-[g_{y}]_{\mathcal{A}},$
$\hat{Q}_{yy}=Q_{yy}+[2\pi_{\rm{I}}-\lambda_{k}]_{+}[g_{yy}]_{\mathcal{A}}+2[g_{y}]^{\mathsf{T}}_{\mathcal{A}}P[g_{y}]_{\mathcal{A}},$ (52)
@@
and projection in (20). $[I_{w}]_{\mathcal{A}}$ is a modified identity matrix whose $i$ th diagonal element is 1 if $g_{i}$ is active (positive after projection in (20)) and zero otherwise. First-order optimality condition for quadratic approximation of $\hat{Q}$ gives
@@ 4.4.2 Multiple-shooting
$\hat{Q}_{\tilde{y}}=l_{\tilde{y}}+h_{\tilde{y}}^{\mathsf{T}}[2\pi_{\rm{E}}-\nu_{k}]+[g_{\tilde{y}}]_{\mathcal{A}}^{\mathsf{T}}[2\pi_{\rm{I}}-\lambda_{k}]_{+},$
$\hat{Q}_{\lambda}=-\frac{1}{2}P_{\rm{I}}^{-1}\big{[}[2\pi_{\rm{I}}-\lambda_{k}]_{+}-\lambda_{k}\big{]},\ \hat{Q}_{\nu}=-P_{\rm{E}}^{-1}[\pi_{\rm{E}}-\nu_{k}],$
@@ 4.5.1 Introducing Copy Variables
Here, we use $Y=[X^{\mathsf{T}},U^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{n+m}$, which is similar to $y_{k}$. This formulation is equivalent to setting $x=[X,U]$, $z=[X^{\mathrm{c}},U^{\mathrm{c}}]$, $A=I,B=-I$, and $c=0$ in (22). The problem’s AL function is given by
@@ 4.5.2 ADMM DDP
ADMM DDP repeats the three update processes in (54), (57), and (58) until the residuals mentioned in section 2.6 become small enough. Upon convergence, the original and copy variables will reach to consensus, and as a result, the final solution will be optimal while satisfying all constraints. It is well known, however, that ADMM might require many iterations until reaching high accuracy Boyd et al. (2011).
@@ 5.2 Multiple shooting-SQP
where the Lagrangian is given by $\mathcal{L}=J+\lambda_{\rm{v}}^{\mathsf{T}}G(X,U)+\nu_{\rm{v}}^{\mathsf{T}}\hat{F}(X,U),$ with Lagrangian multipliers $\lambda_{\rm{v}}\in\mathbb{R}^{wN}$ and $\nu_{\rm{v}}\in\mathbb{R}^{nN}$. After solving the problem above, $Y$ (both $X$ and $U$) is updated by the following linear update law:
@@ 6 Theoretical Time Complexity
In SQPs, the bottleneck is solving the QP subproblems, whose Hessian has the size of $(N-1)m\times(N-1)m$ in single- and $\{(N-1)m+Nn\}\times\{(N-1)m+Nn\}$ in multiple-shooting. These large Hessians lead to $\mathcal{O}(N^{3}m^{3})$ and $\mathcal{O}(N^{3}(m+n)^{3})$ for single- and multiple-shooting, respectively. However, as mentioned in section 5.2, several methods in structure-exploiting multiple-shooting SQP that can reduce the complexity to $\mathcal{O}(N(m+n)^{3})$ are available. In single-shooting, dependency on $N$ can reduced to not linear, but quadratic $\mathcal{O}(N^{2}n^{3})$ (Kouzoupis et al., 2018). From a computational point of view, the single-shooting variant is useful for a problem with a short time horizon $N$ and a large state dimension $n$.
@@ 7 Numerical experiments
    How robust is each algorithm to varying initial conditions and targets?
@@
To answer these three questions, we compare the constrained DDPs and SQPs on four different dynamical systems, that is, an inverted pendulum, a 2D quadrotor with a pendulum (quadpend) based on Singh et al. (2022), a tadpole-like swimmer as in Tassa et al. (2007), and, Franka Emika Panda robotic arm (Haddadin et al., 2022) simulated in Brax with mjx backend Freeman et al. (2021); Zakka et al. (2022). For investigating (R1), we use all systems, while for (R2), we use the quadpend and Panda. In (R3), we focus on the quadpend. The state $x$ of the inverted pendulum consists of the angle and angular velocity of the pendulum as $x=[\theta,\dot{\theta}]^{\mathsf{T}}\in\mathbb{R}^{2}$ and the control $u\in\mathbb{R}$ is the torque applied to the pendulum. The state of the quadpend consists of the 2D position and orientation of the quadrotor, the angle of the pendulum, and the time derivative of them, which leads to $x\in\mathbb{R}^{8}$. We control the force that the two rotors generate, and therefore $u\in\mathbb{R}^{2}$. We choose this example because of the nonlinearity of the constraints by the pendulum part. The swimmer is a more complex system consisting of five links, whose control is torque is applied in the four joints. The dynamics of the swimmer are found in Tassa et al. (2007). The state consists of the 2D position of the nose, four joint angles and their time derivatives. Hence, the state is $x\in\mathbb{R}^{12}$. The control $u\in\mathbb{R}^{4}$ is the torques generated in the joints. The robotic arm Panda has seven joints. The state has the angles of joints and its derivatives, which gives state $x\in\mathbb{R}^{14}$. The control $u\in\mathbb{R}^{7}$ is a command of the joints. The angles and commands have limits. Details of the dynamics of the systems and parameters are provided in Appendix D. All experiments except for Panda were performed with MATLAB (MATLAB, 2022). Experiment of Panda is with JAX (Bradbury et al., 2018).
@@ 7.1.1 Progress of Optimization
as a representative of the gradient of the cost of a trajectory. In a constrained setting, the gradient of $J$ changes to a gradient of the modified objective depending on the algorithm, as shown in the left column of Table 2. For SQP, we use the gradient of the Lagrangian in (26). We note that this is intended for monitoring the optimization of one algorithm, but not for comparing the value across algorithms. We also note that the gradient of the $\log$ barrier DDP has a different property due to the barrier term. In other methods, the gradient may be used as an exit criterion, but in barrier DDP, it is not (Almubarak et al., 2022). Nevertheless, we present it for completeness. We also monitor the values presented in the right column of the table, all of which are related to constraint violation, the residual of optimality condition for dual variables (if the method takes them into account), and penalty parameters (if the method includes any and changes them over iterations). The norm here is the infinity norm taken over all the time steps, e.g.,
@@ 7.1.2 Results
Inverted pendulum: In this task, the goal is to swing up the pendulum while satisfying the constraints $-0.8\leq u_{k}\leq 0.8,-1.5\leq\dot{\theta}\leq 1.5$. The results are provided in Fig. 1. In the right column of each figure, we present the progress of optimization with the gradient of the cost and the other metrics given in Table 2. The vertical dotted lines in AL, PDAL, and ADMM DDP represent different outer loops. In the left column, graphs showing the evolution of $\theta$, $\dot{\theta}$ and control $u$ are illustrated. The constraints and targets are given as dotted lines and circles, respectively. Since constraints are linear in this problem, the Hessian of the constraints is zero. Thus, the exact Hessian and that with the GN approximation are identical. See AL multi exact and AL multi approx. in Fig. (1). $\log$ barrier and IP DDP terminate earlier than other methods because the descent direction cannot be found with a large regularizer in (35). Fig. 2(a) shows the evolution of cost (original cost $J$ in (38)). The largest constraint violation (if any) over iterations is shown in Fig. 2(b). We observe that most methods are able to complete the task with low $(<10^{-5})$ constraint violation. In multiple-shooting DDPs, inequality and equality constraints exhibit comparable magnitudes, contrasting with multiple-shooting SQP, where inequality constraints consistently maintain feasibility. We postulate that this phenomenon arises from the property of SQP, that is, solving QP under constraints, rather than incorporating all elements into the $Q$ function as in DDP. Although both SQP and DDP rely on line search to find a proper step size, SQP seems to better capture the information of constraints. The $\log$ barrier method can keep the trajectory always feasible and therefore does not appear in Fig. 2(b). However, its cost is higher than those of the other DDP methods. This is due to the fixed penalty parameter $\mu$ in (4.1). In exchange for a simple implementation, this fixed $\mu$ makes the algorithm only approximately solve the problem, resulting in a higher cost in this experiment. In this problem, the controller needs to hit its limit over many time steps to complete the task. This makes the problem difficult for ADMM DDP, as we analyzed in section 4.6. In Appendix D.6, we further relax the control limit to $u\in[-0.9,0.9]$ and observe that ADMM DDP is able to handle problems with less tight constraints. For SQP, the multiple shooting version performs similarly to other successful DDP methods, whereas the single shooting variant performs poorly. This difference comes from the fact that the cost function in the multiple-shooting SQP is strictly quadratic in the decision variables (state and control); on the other hand, in single-shooting, the cost is not exactly quadratic in control due to the elimination of state via linearized dynamics.
@@
Quadpend: The quadpend is navigated to reach the target $x_{\rm{g}}=[2.5,-1,0,{\pi}/{2},0,0,0,0]^{\mathsf{T}}$ (pendulum upright) from the initial state $x_{0}=[-2,1,0,0,0,0,0,0]^{\mathsf{T}}$ (pendulum down) while avoiding four obstacles under box control constraints. The initial trajectory is shown in Fig. LABEL:fig:quad_init_last, which is obtained with single-shooting PDAL DDP. We show the results of single and multiple shooting methods in Fig. 3. In this experiment, multiple-shooting AL DDPs reached maximum iteration by reaching a plateau where the inner loops could not find a descent direction. The multiple-shooting SQP algorithm terminated due to failure in the line search to find a descent direction even after regularizing the Hessian as explained in Section 5.1. Here, the GN approximation makes a difference because of nonlinear constraints. Indeed, we observe the effectiveness of the approximation as multiple-shooting AL DDP fails without the approximation, getting stuck at a poor local minimum with a large constraint violation. Other than the multiple-shooting AL DDP with exact Hessian, all algorithms can let the quadpend hit the target with the pendulum up, although the solutions vary. This is because the algorithms are local methods that use the local approximation of the cost and dynamics.